a rope measures 168 and is cut in three peices and the first peice is one-half bigger than the second peice and the third peice is three time the second peice
Answer by mananth(16946)

let second piece be x
first piece = 1.5x
third piece =3x
...
add up
x+1.5x+3x=168
5.5x=168
11/2x=168
11x=2*168
11x=338
x= 338/11
x=30.7
3x=92.1
1.5x=46.05
...
m.ananth@hotmail.ca